Simplify the equation :(cosec a+cot a)(1-cos a).
i want the answer immediately please......

Grade:10

1 Answers

Shreyash Rane
7 years ago
(cosec a+cot a)(1-cos a)
= cosec a-cosec a*cos a+cot a-cot a*cos a
=1/sin a-(cos a)/sin a+(cos a-cos^2 a)/sin a
=(1-cos a+cos a-cos^2 a)/sin a
=(1-cos^2 a)/sin a
=sin^2 a/sin a = sin a